A fruit seller buys 50 bananas at the market price of 40 bananas from market,if he sells these bananas giving a discount of 5% ,what is the profit % ?
1
Expert's answer
2014-12-08T11:08:07-0500
Q = 50 bananas, P = $40.If he sells these bananas giving a discount of 5%, the profit is TP = TR - TC = P*Q - TC = 50*40*(1 - 0.95) = $100.